Minutes of the forward endcap EMC eZuce SRN meeting, 18th of July 2016, 14:00-15:10

Presence:

  • Basel: Werner
  • Bochum: Cathrina, Claudius, Fritz-Herbert, Malte, Matthias, Miriam, Patrick, Stephan, Tobias H., Tobias T., Tom
  • Bonn: Christoph
  • Giessen: Hans-Georg
  • Groningen: Myroslav
  • Stockholm: -
  • Uppsala: Pawel
  • GSI: -

APD subunits:

  • Summary of new insights into the '1:2 problem' (preamp output ratio of 2 of the two APDs of one crystal, seen on half of the crystals)
    • The unit showing both ratios of 1:1 and 1:2 in one cosmics run was found to be completely decoupled from the crystal
    • Looking at the event numer (aka time) one sees that the ratio started to be 1:2 at the beginning and soon turned to 1:1
    • A 'pressed coupling' of the APD unit against the crystal with the remaining silicon pad still attached to the APDs gives a read out ration of 1:1
  • It seems that the 1:2 ratio problem is a coupling problem even though we are far from fully understanding what is going on
    • Does a perfect adhesive coupling give twice the light output of a coupling with the silicon just pressed to the crystal? (Measurements explicitly adressing this question are going on)
    • Why do we see ratios of 1:2 directly after a fresh glueing?
    • Why does it seem there are only tight coupled or totally decoupled APDs (no ratios in betwwen 1:1 and 1:2)? (There seem to be no 'partly' (de)coupled APDs)

APD preamps:

  • There are about 170 VPTT preamps of one gain ready and tested that will be sent this week or next week
  • There are also 32 APD preamps of the latest version ready to be sent

Radiation map:

  • There was some discussion on how to proceed with getting detailed, realistic radiation maps
  • It was agreed that Karoly, who started working on this, should go on with work on this topic
  • Pawel has been in contact with Karoly as he needs information for further tests, awaits approximate values this week

ADC boards:

  • Pawel needs information about the measured ratio between high and low gain in order to modify for a ratio of 1:16 (Malte will provide the Bochum results)
  • We decided to leave all four different shaping constants implemented to the boards for the time being as we need to closer look at pulse shape stability
    • Myroslav pointed out that pulse shape stability is an important issue in the deconvolution of pile up events as the algorithms rely on stable shapes

HV adjustment boards:

  • Christoph showed pictures of the latest PCB layouts
  • There are mock-ups in the works to mechanically test for connectability, mounting sequence etc.
  • For I2C cabling the barrel 'Bedea' cable seems to be probably not suitable (too tiny), Christoph will ask for a suitable cable w/ two differential lines plus one power line in a 3...4 mm overall diameter cable

HV module repair:

  • We will send three broken HV modules to ISEG and await offers for repair.
  • We will continue testing until all HV modules currently in Bochum have been screened

Irradiation of 'inner endcap wall' material candidates:

  • Hans reported of 'plastic smell' at the end of an exposure to 1000 Gy of four different foam test pieces
  • It is unclear what piece caused the smell
  • He will send back the pieces to Bochum for evaluation of material deterioration
  • Once we decided for a material we will need to test it at twice the dose as 1000 Gy may be a too short estimate for the inner walls
-- ThomasHeld - 21 Jul 2016

Added details regarding HV module repair -- TobiasTriffterer - 21 Jul 2016
Topic revision: r2 - 21 Jul 2016, TobiasTriffterer
 
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding PANDA Wiki? Send feedback
Imprint (in German)
Privacy Policy (in German)